OpenStudy (anonymous):

do with the quadratic formula or equation. x^2 + 4x + 4 = 7????

OpenStudy (radar):

That sets up nicely for factoring

OpenStudy (anonymous):

I would always try to factor first just because its faster personaly i only use the quadratic if I have to

OpenStudy (anonymous):

my teacher told me to put it into a=1 b=4 c= - 3 form or something and then put it into the equation (i know what the equation is im just stuck)

OpenStudy (anonymous):

ok then it sounds like you are going to have to do it with the quadratic.

OpenStudy (radar):

I agree, I didn't see the 7

OpenStudy (anonymous):

subtract both sides by 7 then you get x^2=4x-3 a=1 b=4 c =-3 use qaud eq x=(-b+/-sqrt(b^2-4ac))/(2a)

OpenStudy (anonymous):

i keep coming up with -2 + or - radical 7

OpenStudy (anonymous):

is that wrong or right im confused

OpenStudy (radar):

That looks like it would do it.

OpenStudy (anonymous):

did you apply the formula correctly? If you did, you should get the right answer.

OpenStudy (anonymous):

o okay thank u so much!

OpenStudy (anonymous):

should get (-4+sqrt28)/2 and (-4-sqrt28)/2

OpenStudy (radar):

\[-2\pm \sqrt{7}\]

OpenStudy (anonymous):

i got what radar said. not eric8 ahh which one is right

OpenStudy (radar):

I think EricC8 can be simplified for example the 28 is 7 x 4 and the sq root of 4 could be taken out of the radical

OpenStudy (anonymous):

o okay thank u so much!

OpenStudy (radar):

then the denominator 2 could be divided into each term

OpenStudy (anonymous):

oh i see what you mean

OpenStudy (radar):

By Jove I believe you've got it!

OpenStudy (anonymous):

haha thank u so much now ill be able to get an a+ on the test (:

OpenStudy (radar):

Hope so, good luck on the test.

OpenStudy (anonymous):

thanks